TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint5x2006x10x02x11x17

hwc

Please find below the plot for RR53, today at 13:21 hours. The system is working according to spec. (high speed since about 10:00 h). Keep in mind that the nominal values requested by specification are: Temp (mean) ground floor: 28 (+- 5) C Temp. (mean) top floor: 30 (+- 5) C

The measured values are (at 14:04 h) Temp. (mean) GF: 24.2 C Temp (mean) top : 30 C (vault) Temp mean regulation: 27.1 C

Water monitoring in UW45 is also underway and according to spec. Plot in the coming minutes.

We are at this moment discussing with J.C. Guillaume (on site) some adjustmemts to try and improve the temperature of his cables.

JPEG Outlook.jpg:
Outlook.jpg

Comments:

Topic attachments
I Attachment History Action Size Date Who Comment
JPEGjpg Outlook.jpg r1 manage 142.9 K 2006-10-02 - 11:19 JacekSzkutnik  
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-10-02 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login